Angelo’s marks on the first three tests are 10/20, 15/20, and 17/20 what mark would he have to get on the next test to make his average on all four tests 15/20 ?
a) 15/20 b) 16/20 c) 17/20 d) 18/20 e) 20/20

Answer by edjones(7311)

(10/20 + 15/20 + 17/20 + x)/4=15/20
x + 42/20 = 60/20
x=60/20 - 42/20
= 18/20
.
Ed